# How many simulations do we need for simulation-based inference in cosmology?

### v1

@@
Abstract
How many simulations do we need to train machine learning methods to extract information available from summary statistics of the cosmological density field? Neural methods have shown the potential to extract non-linear information available from cosmological data. Success depends critically on having sufficient simulations for training the networks and appropriate network architectures. In the first detailed convergence study of neural network training for cosmological inference, we show that currently available simulation suites, such as the Quijote Latin Hypercube(LH) with 2000 simulations, do not provide sufficient training data for a generic neural network to reach the optimal regime, even for the dark matter power spectrum, and in an idealized case. We discover an empirical neural scaling law that predicts how much information a neural network can extract from a highly informative summary statistic, the dark matter power spectrum, as a function of the number of simulations used to train the network, for a wide range of architectures and hyperparameters. We combine this result with the Cramer-Rao information bound to forecast the number of training simulations needed for near-optimal information extraction. To verify our method we created the largest publicly released simulation data set in cosmology, the Big Sobol Sequence(BSQ), consisting of 32,768 $\Lambda$CDM n-body simulations uniformly covering the $\Lambda$CDM parameter space. Our method enables efficient planning of simulation campaigns for machine learning applications in cosmology, while the BSQ dataset provides an unprecedented resource for studying the convergence behavior of neural networks in cosmological parameter inference. Our results suggest that new large simulation suites or new training approaches will be necessary to achieve information-optimal parameter inference from non-linear simulations.
preprint: APS/123-QED
I Introduction
The standard model of cosmology[1, 2] has emerged from compelling observational evidence across multiple scales and probes. Dark matter, first indicated by galaxy rotation curves[3, 4, 5] and later confirmed through weak gravitational lensing[6], comprises approximately 25% of the universe’s energy content, as measured through cosmic microwave background (CMB) observations from WMAP and Planck[7, 8]. CMB data reveals that ordinary matter and dark matter together account for only about 30% of the critical density needed for a flat universe. This finding, combined with the discovery of cosmic acceleration[9, 10], points to the existence of dark energy—represented in $\Lambda$CDM by a cosmological constant ($\Lambda$)—which makes up the remaining 70% of the cosmic energy budget.
Thus, a major effort in modern cosmology is directed towards inferring the cosmological parameters from astronomical observations and testing the fundamental tenets of $\Lambda$CDM. This endeavor is supported by increasingly precise numerical simulations[11] and forthcoming data from several major cosmological surveys, including Euclid[12], DESI[13], Vera Rubin Observatory[14], the Simons Observatory[15], and CMB-S4[16].
@@ I Introduction
While sophisticated methods for cosmological parameter inference have dramatically reduced our uncertainty about cosmological parameters in recent decades[17, 18, 19, 20, 21, 22], current and upcoming surveys present new methodological challenges. The traditional approach compares summary statistics—such as n-point correlation functions, number counts, and wavelet statistics—against theoretical predictions or numerical simulations[23, 24, 25, 26]. These methods typically assume an explicit form of the likelihood function, often Gaussian, but this assumption becomes problematic in the non-linear regime where the true likelihood form may not be known, especially in the regime of a few modes and complex modeling. For example, even when the number of modes is large, non-linear evolution can lead to strong correlations, slowing the asymptotic approach to the Gaussian form. Even when good approximations do exist, standard Bayesian inference techniques like Markov Chain Monte Carlo (MCMC) can face convergence challenges[27].
@@
Implicit Inference (II), or Simulation-Based Inference (SBI), has emerged as a promising alternative that learns the joint statistics of data and parameters directly from simulated pairs, avoiding the need for explicit likelihood assumptions. SBI has demonstrated both accuracy and computational efficiency across various applications[28, 29, 30, 31, 32, 33, 34]. Neural networks play two crucial roles in modern SBI approaches: generating or compressing informative summary statistics and representing statistical distributions of likelihood or posterior. Information Maximizing Neural Networks (IMNN) exemplify the first role, focusing on Fisher information maximization[35], while (ensembles of) normalizing flows[36, 37] and diffusion models[38, 39, 40] excel at the second.
@@
Given the centrality of trained neural networks to this powerful approach, a key question emerges: How many simulations are needed to achieve optimal results? While one could generate simulations until the extracted information plateaus, this brute-force approach is computationally expensive or even intractable. Our work presents the first comprehensive analysis of neural network convergence in cosmological information extraction and provides a predictive framework to determine the required number of simulations.
@@
For our investigation, we focus on the power spectrum $P(k)$ as our summary statistic. This choice is motivated by the fundamental role of $P(k)$ in cosmology: at large scales and high redshifts, where the universe remains approximately Gaussian and homogeneous, $P(k)$ captures the complete statistical description[41, 27, 42, 43, 24]. While current three-dimensional surveys access significant non-Gaussian information at smaller scales and lower redshifts, any method claiming to extract cosmological information should, at minimum, demonstrate proficiency with $P(k)$ in the non-linear regime. Our findings reveal that even for this basic case, the required training set size is substantial, suggesting that simulation requirements may limit the amount of information extracted from cosmological data and will be a significant challenge for brute-force neural network-based analyses.
@@
The remainder of this paper is organized as follows. Section II describes our simulation suite. Section III presents methods for measuring parameter information content in cosmological datasets. Section IV details our neural network methodology. In section V we present our method for estimating the number of required training sets to asymptote to the Cramer Rao information bound. In order to test these predictions we present the Big Sobol Sequence (BSQ) in section V.1 comprising 32,768 simulations designed specifically for testing neural network convergence and quantifying information content. Our analysis includes careful consideration of the important topic of network architecture selection (VIII.1) and hyperparameter optimization (VIII.1). We discuss our findings and conclude in section VI.
II The Quijote Simulation Suite
The Quijote Simulation Suite is a collection of more than 88,000 full N-body dark matter simulations designed to probe the large-scale structure of the universe[11]. Using the TreePM code GADGET-III which is an improved version of GADGET-II[44], these simulations span more than 40,000 cosmological models, exploring a parameter space defined by ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8},M_{\nu},w}$. Each simulation models the evolution of dark matter in a cubic volume of 1$(Gpc/h)^{3}$, computed at three spatial resolutions: $256^{3}$, $512^{3}$, and $1024^{3}$ grid points.
@@ II The Quijote Simulation Suite
The suite includes several complementary subsets designed for different analysis approaches. For instance, the Latin Hypercube (LH) subset consists of 2,000 simulations sampling five parameters ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}}$, specifically configured for machine learning applications. These parameters are sampled uniformly within ranges centered around the Planck 2018 best-fit values[2]. Additionally, Quijote provides 15,000 fiducial simulations and 500 pairs of finite difference simulations for robust statistical analysis.For our investigation of simulation-based inference using neural networks, we focus on the matter power spectrum $P(k)$ measured at redshift $z=0$. We analyze scales up to $k_{max}=0.804,h/\text{Mpc}$, corresponding to the $0.25$ times Nyquist frequency of a $1024^{3}$ grid in our 1$(Gpc/h)^{3}$ volume. The power spectrum is evaluated at multiples of the fundamental frequency, yielding a 128-dimensional measurement vector over this $k$ range.
III Information Content: Fisher Information and Forecasts
@@
The Fisher information formalism [45, 28] provides a framework for quantifying the information about cosmological parameters $\theta$ contained in data $d(\theta)$. This approach yields insights into maximum likelihood parameter estimates and their uncertainties. For unbiased estimators, the Cramér-Rao information inequality establishes a fundamental lower bound on parameter variance
@@
where $F$ is the Fisher information matrix, defined as: $F=-\langle\nabla_{\theta}\nabla^{T}_{\theta}\mathcal{L}\rangle=\langle\nabla_{\theta}\mathcal{L}\nabla^{T}_{\theta}\mathcal{L}\rangle$, and $\mathcal{L}$ is the log-likelihood of the statistic. Here, the expectation value is taken over different realizations at fixed $\theta$, and the Fisher matrix in equation (1) is evaluated at the fiducial parameter value $\theta=\theta_{\star}$.
In the asymptotic limit of a large number of modes, the power spectrum likelihood is approximately Gaussian, so the log-likelihood takes the form
@@
leading to a Fisher information matrix[46] of
@@
The convergence of both $\nabla_{\theta}\mu$ and the covariance matrix $C$ in equation (3) depends critically on the number of simulations. The Quijote suite provides sufficient simulations for convergence, with 500 pairs of finite difference simulations and 15,000 fiducial simulations.
@@
$\nabla_{\theta}\mathcal{L}|_{\textbf{ML}}=\nabla{\theta}\mu^{T}|_{\textbf{ML}}C^{-1}(d-\mu(\theta_{\textbf{ML}}))=0.$
@@
$\theta_{\textbf{ML}}=\theta_{\star}+F^{-1}\nabla_{\theta}\mu^{T}({\theta_{\star}})C^{-1}(d-\mu(\theta_{\star}))$ (4)
at leading order in $\theta_{\textbf{ML}}-\theta_{\star}$. For parameters far from the fiducial point, equation (4) must be iterated until convergence, requiring covariance matrix evaluation at each intermediate point. However, this iteration is computationally prohibitive due to the large number of simulations required at each step.
@@ IV Neural Estimator
IV.1 Selection of the neural architecture
Multi-layer Perceptrons (MLPs) demonstrate superior performance compared to linear regression models for nonlinear parameter estimation tasks. As fully connected feed-forward neural networks with multiple hidden layers and nonlinear activation functions, MLPs excel at capturing complex relationships in data. We employ an MLP to infer cosmological parameters from simulated power spectra $P(k)$. After extensive architecture exploration, we developed a model (Figure 1) consisting of 6 hidden layers, each followed by LeakyReLU(0.5) activation. The hidden layer dimensions follow an expanding-contracting pattern $(128\rightarrow 512\rightarrow 1024\rightarrow 1024\rightarrow 512\rightarrow 128)$, which improves parameter estimation performance. We will study the effect of variations in the architecture in subsection VIII.1.
@@ IV.1 Selection of the neural architecture
$\mathcal{L}$ $=\sum_{i}\ln(MSE)_{i}$
$=\sum_{i}\ln\left(\frac{1}{N_{batch}}\sum_{n=0}^{N_{batch}}\big{[}\theta_{i}^{(n)}-\hat{\theta}_{i}^{(n)}\big{]}^{2}\right)$ (5)
where $i$ denotes different parameter classes. The learning rate is reduced by a factor $\gamma=0.9$ every 10 epochs using PyTorch’s StepLR scheduler. These hyperparameters have been chosen after an extensive hyperparameter search, as shown in VIII.1. Training the model on the 2000 Latin Hypercube $P(k)$ samples for 500 epochs required approximately one hour on a V100 GPU, with model weights at the lowest validation loss retained for inference.
@@
V Convergence Analysis and Scaling to estimate simulation requirements
@@
Figure 2: Loss ((IV.1)) on held out test data as a function of number of Quijote LH simulations. The loss asymptotes to the Cramér-Rao (C-R) bound $\mathcal{L}_{CR}$ by power law decay $\propto N^{-0.39}$ (cf. (7)). The horizontal line marks where the log loss deviates from the C-R bound by O(1). Based on our initial training set of 1500 we predict that the information extracted by the neural summary is not yet optimal and that nearly 4000 simulations are needed to be in the asymptotic regime. This prediction is verified in Fig. 5.
The figure reveals that 1500 simulations are insufficient to reach the optimal regime. While determining the precise scaling relationship from limited samples is challenging, a simple linear extrapolation in the log-log plot suggests approximately 4000 simulations are needed to saturate the information content of $P(k)$. Drawing insights from recent scaling studies of transformer models[47, 48], we hypothesize that the relationship follows a power law that asymptotes with the increasing number of simulations $N$
@@
where $c$ is a constant and power law fitting to the test losses yields $\alpha\approx 0.39$. Although decreasing log MSE loss does not guarantee increased Fisher information across all parameters, we expect information gains for at least some parameters. In the ideal case, where loss contributions from each parameter are comparable, uncertainties should decrease uniformly across all parameters until information saturation. Figure 3 demonstrates the fraction of optimal information extracted by neural summaries at the fiducial parameter point, when trained on different numbers of simulations. More specifically, the parameter uncertainties are derived from the Fisher information matrix based on the neural summaries, calculated using (3).
@@
V.1 The Big Sobol Sequence (BSQ) Simulations
Our analysis of the Quijote Latin Hypercube simulations indicates the need for a larger training set to saturate the available cosmological information. To verify our predictions about the required simulation count and test convergence behavior, we generated a new suite of 32,768 N-body simulations. These simulations sample the parameter space ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}}$ using a Sobol sequence design. We utilize 16,600 of these simulations, maintaining the same 15:3:2 ratio for train, validation, and test sets as in our previous analysis. To ensure direct comparability, we employ the identical PowerSpectraNet architecture and configuration used for the Latin Hypercube simulations.

### v2

# The Big Sobol Sequence: how many simulations do we need for simulation-based inference in cosmology?

@@
11institutetext: CNRS & Sorbonne Université, Institut d’Astrophysique de Paris (IAP), UMR 7095, 98 bis bd Arago, F-75014 Paris, France
11email: bairagi@iap.fr
22institutetext: Department of Physics and Astronomy, Johns Hopkins University, 3400 North Charles Street, Baltimore, MD 21218, USA 33institutetext: Department of Applied Mathematics and Statistics, Johns Hopkins University, 3400 North Charles Street, Baltimore, MD 21218, USA 44institutetext: Center for Computational Astrophysics, Flatiron Institute, 162 5th Avenue, New York, NY 10010, USA 55institutetext: Department of Astrophysical Sciences, Princeton University, Peyton Hall, Princeton, NJ 08544-0010, USA
The Big Sobol Sequence: how many simulations do we need for simulation-based inference in cosmology?
Anirban Bairagi 11 0009-0009-3089-052X    Benjamin Wandelt 11223344 0000-0002-5854-8269    Francisco Villaescusa-Navarro 4455 0000-0002-4816-0455
[…]
Key Words.:
Large Scale Structure, Simulation based inference, Neural Networks
1 Introduction
[…]
@@ 1 Introduction
While sophisticated methods for cosmological parameter inference have dramatically reduced our uncertainty about cosmological parameters from past surveys, e.g., WMAP (Bennett et al. 2003), Planck (Planck Collaboration et al. 2020), DES (Abbott et al. 2022), BOSS (Dawson et al. 2013; Alam et al. 2017), DESI (DESI Collaboration et al. 2024) in recent decades, current and upcoming surveys present new methodological challenges. The traditional approach compares summary statistics (Peebles 1993, 1980)—such as n-point correlation functions (Peebles 2001; Scoccimarro 2000; Philcox & Slepian 2022), number counts (Artis et al. 2025) from various probes, i.e., galaxy clustering (Groth et al. 1977; Kaiser 1987) and weak lensing (Heymans et al. 2013, 2021; Liu et al. 2015)—against theoretical predictions or numerical simulations. These methods typically assume an explicit form of the likelihood function, often Gaussian, but this assumption becomes problematic in the non-linear regime where the true likelihood form may not be known, especially in the regime of a few modes and complex modeling. For example, even when the number of modes is large, non-linear evolution can lead to strong correlations, slowing the asymptotic approach to the Gaussian form (Sellentin & Heavens 2017; Hahn et al. 2019). Even when good approximations do exist, standard Bayesian inference techniques like Markov Chain Monte Carlo (MCMC) can face convergence challenges (Liddle 2009; Leclercq et al. 2014).
@@
Implicit Inference (II), or Simulation-Based Inference (SBI), has emerged as a promising alternative that learns the joint statistics of data and parameters directly from simulated pairs, avoiding the need for explicit likelihood assumptions. SBI has demonstrated both accuracy and computational efficiency across various applications (Alsing et al. 2018, 2019; Cranmer et al. 2020; Jeffrey & Wandelt 2020; Joachimi, B. et al. 2021; Cole et al. 2022; Lemos et al. 2023). Neural networks play two crucial roles in modern SBI approaches: generating or compressing informative summary statistics (Jeffrey & Wandelt 2020; Lemos et al. 2024; Novaes et al. 2025) and representing statistical distributions such as the likelihood or posterior. Information Maximizing Neural Networks (IMNN) exemplify the first role, focusing on Fisher information maximization (Charnock et al. 2018), while (ensembles of) normalizing flows (Papamakarios et al. 2019, 2017), neural ratio estimation (Cole et al. 2022), and diffusion models (Song et al. 2020b; Ho et al. 2020; Song et al. 2020a) are examples of the second.
While both of these tasks will require assessment of convergence, in this paper we will focus on the question of information extraction [^1] For a recent paper addressing the convergence of neural density estimators in the cosmological context, see Homer et al. (2025)., for the following reasons: 1) suboptimal information extraction will affect downstream tasks such as the inference of the likelihood or the posterior distribution; and 2) it is well known that the difficulty of fitting general high-dimensional distributions scales poorly with the number of dimensions, highlighting the utility of extracting information in a low-dimensional feature vector. Even in cases where neural density estimators are trained conditional on the feature vector, it has been observed in several of the above-cited works that providing long, uncompressed data vectors to conditional neural density estimators leads to suboptimal performance.
@@
Given the centrality of trained neural networks for the extraction of informative features, a key question emerges: How many simulations are needed to achieve near-optimal results? While one could generate simulations until the extracted information plateaus, this brute-force approach is computationally expensive or even intractable. While even suboptimal information extraction may be an improvement over traditional summary statistics, it is clearly interesting to have a simple tool based on a power law fit to assess whether the extracted information from valuable data could be improved by running more training simulations. Our work presents the first comprehensive analysis of neural network convergence in cosmological information extraction and for relatively low-dimensional summary statistics such as the power spectrum provides a predictive framework to determine the required number of simulations based on numerical evaluation of the Fisher information.
@@
For our investigation, we focus on the non-linear dark matter power spectrum $P(k)$ as our main summary statistic. This choice is motivated by the fundamental role of $P(k)$ in cosmology: at large scales and high redshifts, where the universe remains approximately Gaussian and homogeneous, $P(k)$ captures the complete statistical description (Kaiser 1987; Peacock & Dodds 1996; Gil-Marín et al. 2012; Vlah et al. 2016). While current three-dimensional surveys access significant non-Gaussian information at smaller scales and lower redshifts, any method claiming to extract cosmological information from the non-linear regime should, at minimum, demonstrate the ability to do so with the power spectrum (Tegmark et al. 2004; Seljak et al. 2005; Philcox & Ivanov 2022; D’Amico et al. 2024). Our findings reveal that even for this basic case, the required training set size is substantial, suggesting that simulation requirements may limit the amount of information extracted from cosmological data and will be a significant challenge for brute-force neural network-based analyses.
To test whether the notion of scaling laws extends to summary statistics beyond the power spectrum, and other loss functions, we briefly investigate the problem of neural posterior estimation with Kullback-Leibler loss, based on wavelet scattering coefficients.
@@
The remainder of this paper is organized as follows. Section 2 describes our simulation suite. Section 3 presents methods for measuring parameter information content in cosmological datasets. Section 4 details our neural network methodology. In section 5 we present our method for estimating the number of required training sets to asymptote to the Cramér-Rao information bound. In order to test the extrapolation of the resulting power law scaling predictions we present the Big Sobol Sequence (BSQ) in section 5.1 comprising 32,768 simulations designed specifically for testing neural network convergence and quantifying information content. Our analysis includes careful consideration of the important topic of network architecture selection and hyperparameter optimization (see Appendix A). Testing whether the finding of power law scaling translates to other summary statistics and loss functions, we find in Section 5.2 a scaling law for neural posterior estimation based on wavelet scattering transform coefficients with Kullback-Leibler loss function. We discuss our findings and conclude in section 6.
2 The Quijote simulation suite
[…]
@@ 2 The Quijote simulation suite
The suite includes several complementary subsets designed for different analysis approaches. For instance, the Latin Hypercube (LH) subset consists of 2,000 simulations sampling five parameters $\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}$}, specifically configured for machine learning applications. These parameters are sampled uniformly within ranges centered around the Planck 2018 best-fit values (Planck Collaboration et al. 2020).
Additionally, the Quijote suite provides a large set of simulations designed to facilitate the computation of Fisher matrices for arbitrary, informative statistics which we will discuss in section 3. This set contains 15,000 simulations at a fixed fiducial cosmology of $\Omega_{m}=0.3175,$ $\Omega_{b}=0.049,$ $h=0.6711,$ $n_{s}=0.96244$, and $\sigma_{8}=0.834$ and 500 pairs of simulations to support finite difference computations. These pairs are symmetrically displaced from the fiducial cosmology along each parameter direction. Each pair of simulations has identical initial conditions, suppressing sample variance in the finite difference computation.
[…]
3 Information content: Fisher information and forecasts
@@
The Fisher information formalism (Tegmark et al. 1997; Tegmark 1997; Alsing & Wandelt 2018) provides a framework for quantifying the information about cosmological parameters $\theta$ contained in data $d(\theta)$. This approach yields insights into maximum likelihood parameter estimates and their uncertainties. For unbiased estimators, the Cramér-Rao information inequality (Rao 1945; Cramér 1946) establishes a fundamental lower bound on parameter variance
@@
where $F$ is the Fisher information matrix, defined as: $F=-\langle\nabla_{\theta}\nabla^{T}_{\theta}\mathcal{L}\rangle=\langle\nabla_{\theta}\mathcal{L}\nabla^{T}_{\theta}\mathcal{L}\rangle$, and $\mathcal{L}$ is the log-likelihood of the statistic. Here, the expectation value is taken over different realizations at fixed $\theta$, and the Fisher matrix in Eq. (1) is evaluated at the fiducial parameter value $\theta=\theta_{\star}$.
In the asymptotic limit of a large number of modes, and for informative statistics, the power spectrum likelihood is approximately Gaussian, so the log-likelihood takes the form
@@
leading to a Fisher information matrix (Alsing & Wandelt 2018) of
@@
The convergence of both $\nabla_{\theta}\mu$ and the covariance matrix $C$ in Eq. (3) depends critically on the number of simulations Coulton & Wandelt (2023). Since the dark matter clustering statistics used in this paper are highly informative, we have found that the number of simulations provided in the Quijote suite (see Section 2) suffice for computing the finite difference approximation of the derivatives in Eq. (3).
@@
$\nabla_{\theta}\mathcal{L}|_{\text{ML}}=\nabla{\theta}\mu^{T}|_{\text{ML}}C^{-1}(d-\mu(\theta_{\text{ML}}))=0.$
@@
$\theta_{\text{ML}}=\theta_{\star}+F^{-1}\nabla_{\theta}\mu^{T}({\theta_{\star}})C^{-1}(d-\mu(\theta_{\star}))$ (4)
at leading order in $\theta_{\text{ML}}-\theta_{\star}$. For parameters far from the fiducial point, Eq. (4) must be iterated until convergence, requiring covariance matrix evaluation at each intermediate point. However, this iteration is computationally prohibitive due to the large number of simulations required at each step.
@@ 4 Neural estimator
Selection of the neural architecture
Multi-layer Perceptrons (MLPs) demonstrate superior performance compared to linear regression models for nonlinear parameter estimation tasks. As fully connected feed-forward neural networks with multiple hidden layers and nonlinear activation functions, MLPs excel at capturing complex relationships in data. We employ an MLP to infer cosmological parameters from simulated power spectra $P(k)$. After extensive architecture exploration detailed in Appendix A, we developed a model named PowerSpectraNet (Figure 1) consisting of 6 hidden layers, each followed by LeakyReLU(0.5) activation. The hidden layer dimensions follow an expanding-contracting pattern $(128\rightarrow 512\rightarrow 1024\rightarrow 1024\rightarrow 512\rightarrow 128)$, which improves parameter estimation performance. We will study the effect of variations in the architecture in Appendix A.
@@
$\mathcal{L}$ $=\sum_{i}\ln\left(\frac{1}{N_{batch}}\sum_{n=0}^{N_{batch}}\big{[}\theta_{i}^{(n)}-\hat{\theta}_{i}^{(n)}\big{]}^{2}\right)$ (5)
where $i$ denotes different parameter classes. The learning rate is reduced by a factor $\gamma=0.9$ every 10 epochs using PyTorch’s StepLR scheduler. These hyperparameters have been chosen after an extensive hyperparameter search, as shown in Appendix A. Training the model on the 2000 Latin Hypercube $P(k)$ samples for 500 epochs required approximately one hour on a V100 GPU, with model weights at the lowest validation loss retained for inference.
@@
5 Convergence analysis and scaling law to estimate simulation requirements
@@
Figure 2: Loss (Eq. 5) on held out test data as a function of number of Quijote LH simulations. The loss asymptotes to the Cramér-Rao (C-R) bound $\mathcal{L}_{CR}$ by power law decay $\propto N^{-0.39}$ (cf. Eq. 7). We mark on the abcissa the asymptotic regime, defined as where the test loss becomes $e$ times the Cramèr-Rao bound. Based on our initial training set of 1500 we predict that the information extracted by the neural summary is not yet optimal and that several thousand simulations are needed to be in the asymptotic regime. This prediction is verified in Figure 5.
Scaling laws have been found to be ubiquitous in deep learning systems (Hestness et al. 2017) including in recent, seminal scaling studies of transformer models (Kaplan et al. 2020; Edwards et al. 2024). Inspired by these studies, and using the fact that the information in finite, noisy data sets is ultimately bounded, we hypothesize that the relationship follows a power law that asymptotes with the increasing number of simulations NN
@@
where $c$ is a constant and the power law fitting to the test losses yields $\alpha\approx 0.39$. The figure reveals that the test log-loss scales as a power law of the number of simulations. The 1500 training simulations available in the Latin Hypercube set are insufficient to reach the asymptotic regime defined here as where the test loss becomes a factor of $e$ times the Cramèr-Rao bound. The power law scaling predicts that several thousand more simulations are needed to saturate the information content of the non-linear matter $P(k)$.
Although decreasing log MSE loss does not guarantee increased Fisher information across all parameters, we expect information gains for at least some parameters. In the ideal case, where loss contributions from each parameter are comparable, uncertainties should decrease uniformly across all parameters until information saturation. Figure 3 demonstrates the fraction of optimal information extracted by neural summaries at the fiducial parameter point, when trained on different numbers of simulations. More specifically, the parameter uncertainties are derived from the Fisher information matrix based on the neural summaries, calculated using (3).
@@
We will now describe the simulation suite we ran and used to test our prediction.
5.1 The Big Sobol Sequence (BSQ) simulations
Our analysis of the Quijote Latin Hypercube simulations indicates the need for a larger training set to saturate the available cosmological information. To verify our predictions about the required simulation count and test convergence behavior, we generated a new suite of 32,768 N-body simulations using Gadget-III. We call these simulations the Big Sobol Sequence (BSQ) and make them publicly available [^2] https://quijote-simulations.readthedocs.io/en/latest/bsq.html. While matched to the original simulations in terms of particle number and volume, the new simulations were run with slightly more stringent force accuracy parameters: PMGRID=2048, ASMTH=3.0, RCUT=6.0, and MaxSizeTimestep=0.005. All other n-body code parameters are identical to the settings in the original Quijote suite.
The simulations cover the same parameter space as the Quijote LH simulations, sampling the five parameters $\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}$} uniformly within ranges centered around the Planck 2018 best-fit values (Planck Collaboration et al. 2020). The BSQ simulations are generated using a Sobol sequence design (Sobol’ 1967) rather than a Latin Hypercube design. We utilize 16,600 of these simulations, maintaining the same 15:3:2 ratio for train, validation, and test sets as in our previous analysis. We have verified that each parameter (and each parameter pair) is sampled uniformly across the parameter range. To ensure comparability, we employ the same PowerSpectraNet architecture shown in Figure 1 and hyperparameters that are identical to the ones used for the Latin Hypercube simulations.